# Can dissipation induce a transition between many-body localized and thermal states?

### v1

@@ Abstract
The many-body mobility edge (MBME) in energy, which separates thermal states from many-body localization (MBL) states, is a critical yet controversial concept in many-body systems. Here we examine the quasiperiodic $t_{1}-t_{2}$ model that features a mobility edge. With the addition of nearest-neighbor interactions, we demonstrate the potential existence of a MBME. Then we investigate the impact of a type of bond dissipation on the many-body system by calculating the steady-state density matrix and analyzing the transport behavior, and demonstrate that dissipation can cause the system to predominantly occupy either the thermal region or the MBL region, irrespective of the initial state. Finally, we discuss the effects of increasing system size. Our results indicate that dissipation can induce transitions between thermal and MBL states, providing a new approach for experimentally determining the existence of the MBME.
@@ I Introduction
Closed quantum many-body systems with disorder or quasiperiodic potentials can exhibit localization review1 ; review2 , wherein the system cannot act as its own heat bath and thus does not reach thermodynamic equilibrium. Many-body localization (MBL) exhibits several intriguing properties, such as the absence of conductivity (even at finite temperatures) review2 ; Basko2006 ; Gornyi2005 , specific spectral properties Huse2007 ; Huse2010 ; Serbyn2016 , entanglement entropy that follows an area law review1 ; review2 , and the slow logarithmic growth of entanglement entropy Chiara2006 ; Prosen2008 ; Moore2012 ; Serbyn2013 . In addition to its fundamental theoretical importance, MBL has potential significant applications in quantum information Altman2015 ; Nayak2014 , time crystals NormanYao , and other areas, which has garnered widespread theoretical and experimental attention. For example, MBL phenomena have been observed in platforms such as cold-atom systems IBloch2015 ; IBloch2016 ; Islam2015 , circuits with superconducting qubits Roushan2017 ; HHWang2021 and trapped-ion JSmith2016 systems. However, even in these highly controlled implementations, MBL systems are influenced by at least a slight coupling to the environment, such as inelastic scattering from lasers, which may ultimately disrupt localization. Furthermore, considering the observation of MBL in traditional solid-state experiments Lake2207 and the applications of MBL phenomena, the impact of dissipation on MBL becomes even more unavoidable. Therefore, the impact of dissipation on MBL has attracted extensive research attention JSmith2016 ; DAHuse2014 ; Johri2015 ; DAHuse2015 ; Hyatt2017 ; Fischer2016 ; Levi2016 ; Medvedyeva ; Everest2017 ; Knap2017 ; LNWu2019 ; JRen2020 ; Gopalakrishnan ; Roeck2017 ; Luitz2017 ; Richter2024 ; Denisov2018 ; Bloch2017 ; Bloch2019 . When coupled to a thermal bath, dissipation can lead to infinite heating in the long-time limit, making MBL typically considered unstable DAHuse2014 ; Johri2015 ; DAHuse2015 ; Hyatt2017 ; Fischer2016 ; Levi2016 ; Medvedyeva ; Everest2017 ; Knap2017 ; LNWu2019 ; JRen2020 ; Gopalakrishnan ; Roeck2017 ; Luitz2017 , as observed in cold-atom experiments Bloch2017 ; Bloch2019 . Although dissipation can disrupt MBL, transport in the dissipation-induced delocalized phase is different from that in a typical thermalized phase Knap2017 . Additionally, once the dissipation is removed, the system returns to a many-body localized state. Can dissipation drive a transition from a MBL state to a thermal state? Conversely, can a certain type of dissipation drive a transition from a thermal state to a MBL state?
@@
The concept of the mobility edge (ME) is a central idea in the field of localization physics. Similar to how MBL can be seen as the counterpart of Anderson localization in many-body systems, a natural question arises: can the concept of the ME be extended to many-body systems? In other words, is there a critical energy that separates thermal states from many-body localized states, known as the many-body mobility edge (MBME)? Numerous numerical results and some experiments suggest the existence of the MBME MBME0 ; MBME1 ; MBME2 ; MBME3 ; MBME4 ; MBME5 ; MBME6 ; MBME7 ; MBME8 . However, due to computational size limitations, the existence of a definitive MBME remains uncertain. For instance, Ref. noME argues that local fluctuations in the system with a putative MBME can act as mobile bubbles, inducing global delocalization, and hence, a MBME cannot exist. If a MBME exists, could the introduction of dissipation lead to some interesting results? Would such results provide insights into determining the MBME?
@@
In this work, we demonstrate that by introducing a special type of controllable dissipation into a system, which, from the perspective of computable sizes, has the MBME, we can drive the system into a steady state, which predominantly consists of either thermal states or many-body localized states, independent of the initial state. Therefore, within the limits of computable finite sizes, we can assert that dissipation can induce transitions between thermal states and MBL states. At the same time, our results also provide a feasible approach for determining the existence of the MBME in simulated systems such as cold atoms.
II Model and Results
@@
When the NN interaction is included, the single-particle ME may lead to the emergence of a MBME. To characterize the localized and thermal properties of this system, we consider the ratio of adjacent energy gaps Huse2007 ; Huse2010 ; Serbyn2016 as $r_{i}=\frac{\min\left(\delta_{i+1},\delta_{i}\right)}{\max\left(\delta_{i+1},\delta_{i}\right)}$, where $\delta_{i}=E_{i}-E_{i+1}$ represents the energy spacing with the eigenvalues $E_{i}$ listed in ascending order. For the system in the thermal region, its level statistics follow the Wigner-Dyson distribution, and the average $\left\langle r\right\rangle$ converges to $0.53$. In the MBL region, the level statistics are Poisson, with $\left\langle r\right\rangle\approx 0.39$. We rescale the many-body spectrum as $\epsilon_{i}=\left(E_{i}-E_{g}\right)/\left(E_{\max}-E_{g}\right)$, where $E_{g}\left(E_{\max}\right)$ is the eigenenergy of the ground (highest excited) state. Based on this, we divide the eigenvalues into $10$ different energy windows and average samples and all gaps in each window to obtain an $\left\langle r\right\rangle$, as shown in Fig. 1(b). We observe that as the strength of the quasiperiodic potential $\lambda$ increases, the states in the uppermost window become localized first, similar to the single-particle case. In the inset of Fig. 1(b), we fix $\lambda=0.8$. The red and blue lines correspond to the $\left\langle r\right\rangle$ of the lowest energy window and the highest energy window, respectively, which converge to $0.53$ and $0.39$. This indicates that the states in the highest window become localized while the states in the lowest window remain thermalized, suggesting the existence of a MBME. In the following discussion, we fix $\lambda=0.8$ and one-third filling.
Figure 1: (a) IPR of various eigenstates plotted against their corresponding eigenvalues $E$ and the strength of the quasiperiodic potential of the $t_{1}-t_{2}$ model. Black dashed line indicates the approximate mobility edge obtained from Eq. (2). (b) Variation of $\left\langle r\right\rangle$ with energy density ($\epsilon$) and quasiperiodic potential strength $\lambda$. Here $t_{1}=1$ and $t_{2}=0.2$, $U=0.9$, the system size $L=18$, the number of particles $N=6$, and the number of samples is $20$, with each sample corresponding to an initial phase $\phi$. Inset: Variation of $\left\langle r\right\rangle$ with system size at one-third filling for the highest (blue) and lowest (red) energy window.
@@
where $a=\pm 1$, and $j=1,2,\ldots,L-1$. This form of dissipation can be implemented using cold atoms in optical superlattices BD1 ; BD2 ; BD3 ; BD4 or through arrays of superconducting microwave resonators BD6 . This bond dissipation maintains the particle number but modifies the relative phase between adjacent lattice sites. It synchronizes them from an in-phase (out-of-phase) mode to an out-of-phase (in-phase) mode when $a$ is set to $-1$ ($1$).
@@
When dissipation is present, we note that even if the steady state of the system resides in a localized region under the basis of the Hamiltonian, the system’s dynamics are not necessarily localized. For instance, dissipation mechanisms like dephasing can disrupt coherence and consequently destroy the system’s localization properties. The bond dissipation described by Eq. (3) in our study expands to include terms related to dephasing dissipation, which also undermine localization. However, in the case of dephasing dissipation disrupting localization, once dissipation is removed, coherence and localization can be restored, completely eliminating the effects caused by dephasing. In contrast, the bond dissipation we discuss selectively stabilizes the system in specific regions through state selection. Upon removing dissipation, the effects it induces do not vanish. This can be observed through the evolution of the density matrix after dissipation is removed. After abruptly removing the dissipation, we observe the evolution of the density matrix: $\rho(t)=\sum_{mn}e^{i(E_{m}-E_{n})t}\rho_{mn}|\psi_{m}\rangle\langle\psi_{n}|$, where $E_{m}$ and $E_{n}$ are the eigenvalues corresponding to the many-body eigenstates $|\psi_{m}\rangle$ and $|\psi_{n}\rangle$. We observe that the diagonal elements of $\rho$ (where $E_{m}=E_{n}$) do not change with time, while the off-diagonal elements fluctuate over time. However, with longer observation times, we need to examine the average effects of the dynamics over this interval, leading to the vanishing of the effects caused by the off-diagonal elements. Therefore, when the steady state primarily occupies the localized (thermalization) region, removing dissipation reveals behavior characteristic of the localized (thermalization) region. Based on this observation, we can conclude that dissipation serves as an intermediate process capable of inducing a transition between thermalization and MBL states.
Experimental platforms such as ultracold atoms transportExp1 ; transportExp2 ; transportExp3 ; transportExp4 ; transportExp5 can probe the transport properties of a system. Next, we will examine the transport properties of the system after it reaches a steady state and the dissipation is removed. By using these properties to distinguish between thermalization and MBL behaviors, we can experimentally verify that dissipation can induce transitions between thermal and localized states. We assume the system reaches a steady state at the time $t_{0}$, after which dissipation is removed. At this point, the system returns to a non-equilibrium state, and we then study its response to a probed electric field. The change in current $\delta I$, for a weak probed electric field, can be derived from linear response theory.
@@
$\sigma(t,t^{{}^{\prime}})=\frac{i}{N}\theta(t-t^{{}^{\prime}})Tr[\rho(t_{0})[\hat{I}(t),\hat{B}(t^{{}^{\prime}})]],$ (6)
where the position operator $B=\sum_{j}jc_{j}^{\dagger}c_{j}$ and the current operator $I=\frac{dB}{dt}=-i[B,H]$. The current change in the thermal region is expected to show large fluctuations because it is significantly affected by the electric field, whereas in the MBL state, the current change should exhibit smaller fluctuations due to minimal influence from the electric field [see the appendix A]. In Fig. 3, the blue line represents the $a=1$ case. As discussed earlier, in this scenario, the steady state primarily occupies the thermalization region near the ground state. After removing dissipation, we observe that the current change consistently exhibits larger oscillations. In contrast, when $a=-1$ (the orange line in Fig. 3), the system’s steady state mainly occupies the many-body localized region near the highest state. After removing dissipation, we see that the current change oscillates with a smaller amplitude and evolves to near zero at a faster rate, exhibiting relatively localized properties.
@@
III Discussion on the effects of system size
As discussed earlier, in small systems, there exists a MBME, and dissipation can induce transitions between thermal and MBL states. However, when the system size increases, the existence of a MBME becomes a controversial issue. If there is no MBME, then the dissipation-induced transitions we discussed here would not exist. We use the density matrix renormalization group (DMRG) method to study the thermalization and localization properties of the ground state and the highest excited state in large systems. To characterize the localization properties of these two states, we consider the density distribution of single-particle excitation, defined as the difference in distribution after introducing an additional particle into the many-body system, i.e., $\delta n_{j}=\rho_{N+1}\left(j\right)-\rho_{N}\left(j\right)$, where $\rho_{N}\left(j\right)=\langle\psi^{N}_{g(e)}|n_{j}|\psi^{N}_{g(e)}\rangle$ and $|\psi^{N}_{g(e)}\rangle$ is the ground (highest excited) state of the system with $N$ particles. We add a particle to a system with size $L=144$ and particle number $N=32$. Fig. 4 shows the distribution of $\delta n$ at each lattice site. We see that for the ground state (the orange dots), the single particle excitation is distributed quite uniformly across different lattice sites, indicating an extended property. In contrast, for the highest excited state (the blue dots), $\delta n_{j}$ is primarily concentrated on a few lattice sites, showing localized behavior. These results correspond to those in Fig. 1, indicating a significant difference in localization properties between the ground state and the highest excited state in larger systems. Similarly, we can also define the IPR of single-particle excitation (SIPR) as

### v2

# Inducing a transition between thermal and many-body localized states and detecting many-body mobility edges through dissipation

@@ Abstract
The many-body mobility edge (MBME) in energy, which separates thermal states from many-body localization (MBL) states, is a critical yet controversial concept in MBL physics. Here we examine the quasiperiodic $t_{1}-t_{2}$ model that features a mobility edge. With the addition of nearest-neighbor interactions, we suggest the potential existence of a MBME. Then we investigate the impact of a type of bond dissipation on the many-body system by calculating the steady-state density matrix and analyzing the transport behavior, and demonstrate that dissipation can cause the system to predominantly occupy either the thermal region or the MBL region, irrespective of the initial state. Finally, we discuss the effects of increasing system size. Our results indicate that dissipation can induce transitions between thermal and MBL states, providing a new approach for experimentally determining the existence of the MBME.
@@ I Introduction
Closed quantum many-body systems with disorder or quasiperiodic potentials can exhibit localization review1 ; review2 ; review3 , wherein the system cannot act as its own heat bath and thus does not reach thermodynamic equilibrium. Many-body localization (MBL) exhibits several intriguing properties, such as the absence of conductivity (even at finite temperatures) review2 ; Basko2006 ; Gornyi2005 , specific spectral properties Huse2007 ; Huse2010 ; Serbyn2016 , entanglement entropy that follows an area law review1 ; review2 , and the slow logarithmic growth of entanglement entropy Chiara2006 ; Prosen2008 ; Moore2012 ; Serbyn2013 . In addition to its fundamental theoretical importance, MBL has potential significant applications in quantum information Altman2015 ; Nayak2014 ; CYin2024 , time crystals NormanYao , and other areas, which has garnered widespread theoretical and experimental attention. For example, MBL phenomena have been observed in platforms such as cold-atom systems IBloch2015 ; IBloch2016 ; Islam2015 , circuits with superconducting qubits Roushan2017 ; HHWang2021 and trapped-ion JSmith2016 systems.
However, even in these highly controlled implementations, MBL systems are influenced by at least a slight coupling to the environment, such as inelastic scattering from lasers, which may ultimately disrupt localization. Furthermore, considering the observation of MBL in traditional solid-state experiments Lake2207 and the applications of MBL phenomena, the impact of dissipation on MBL becomes even more unavoidable. Therefore, the impact of dissipation on MBL has attracted extensive research attention JSmith2016 ; DAHuse2014 ; Johri2015 ; DAHuse2015 ; Hyatt2017 ; Fischer2016 ; Levi2016 ; Medvedyeva ; Everest2017 ; Knap2017 ; LNWu2019 ; JRen2020 ; Gopalakrishnan ; Roeck2017 ; Luitz2017 ; Richter2024 ; Denisov2018 ; Bloch2017 ; Bloch2019 ; Brighi2023 . When coupled to a thermal bath, dissipation can lead to infinite heating in the long-time limit, making MBL typically considered unstable DAHuse2014 ; Johri2015 ; DAHuse2015 ; Hyatt2017 ; Fischer2016 ; Levi2016 ; Medvedyeva ; Everest2017 ; Knap2017 ; LNWu2019 ; JRen2020 ; Gopalakrishnan ; Roeck2017 ; Luitz2017 , as observed in cold-atom experiments Bloch2017 ; Bloch2019 . Although dissipation can disrupt MBL, transport in the dissipation-induced delocalized phase is different from that in a typical thermalized phase Knap2017 .
@@
The concept of the mobility edge (ME) ME1 is a central idea in the field of localization physics. Similar to how MBL can be seen as the counterpart of Anderson localization PWAnderson in many-body systems, the ME can be extended to many-body systems, suggesting the existence of the many-body mobility edge (MBME)– a critical energy in the spectrum that separates thermal and MBL eigenstates. Numerous numerical results and some experiments suggest the existence of the MBME CYin2024 ; MBME0 ; MBME1 ; MBME2 ; MBME3 ; MBME4 ; MBME5 ; MBME6 ; MBME7 ; MBME8 . However, due to computational size limitations, the existence of a definitive MBME remains uncertain. For instance, Ref. noME argues that local fluctuations in the system with a putative MBME can act as mobile bubbles, inducing global delocalization, and hence, a MBME cannot exist. However, this picture has raised some doubts, as Ref. Brighi2020 demonstrates that MBL remains stable in the presence of small bubbles in large dilute systems. If a MBME exists, could the introduction of dissipation lead to some interesting results?
@@
Recently, Liu et al. applied a bond-dissipative operator (given by Eq. (3) below) to a system with single-particle MEs and discovered that it can induce a transition between extended and localized states BD5 . If such dissipation is applied to systems with MBMEs, could it induce a transition between many-body localized and thermal states? If so, this would mean that we can manipulate the transport properties of these many-body systems. Additionally, it would imply that we could separately measure the properties of many-body localized and thermal states, providing new directions for experimental detection of MBMEs. In this work, we demonstrate that by introducing such dissipation into a system that, in terms of its computable sizes, exhibits the MBME, we can drive the system into a steady state, which predominantly consists of either thermal states or many-body localized states, independent of the initial state. Therefore, within the limits of computable finite sizes, we can assert that dissipation can induce transitions between thermal states and MBL states. At the same time, our results also provide a feasible approach for determining the existence of the MBME in simulated systems such as cold atoms.
II Models and Results
II.1 Quasiperiodic $t_{1}-t_{2}$ model
@@
When the NN interaction is included, the single-particle ME may lead to the emergence of a MBME. To characterize the localized and thermal properties of this system, we consider the ratio of adjacent energy gaps Huse2007 ; Huse2010 ; Serbyn2016 as $r_{i}=\frac{\min\left(\delta_{i+1},\delta_{i}\right)}{\max\left(\delta_{i+1},\delta_{i}\right)}$, where $\delta_{i}=E_{i}-E_{i+1}$ represents the energy spacing with the eigenvalues $E_{i}$ listed in ascending order. The range of $r$ is from 00 to $1$, making it more convenient to use $r$ instead of $\delta$ for analysis. $r$ reflects the uniformity of the level spacing distribution. When the level distribution is perfectly uniform, meaning all level spacings are equal, $r=1$. Therefore, a larger $r$ indicates a more uniform level distribution. For the system in the MBL region, the absence of correlation between energy levels results in Poissonian level statistics, with $\left\langle r\right\rangle\approx 0.39$. In the thermal region, its level statistics follow the Wigner-Dyson distribution, where level repulsion leads to a more uniform distribution, and the average $\left\langle r\right\rangle$ converges to $0.53$. We rescale the many-body spectrum as $\epsilon_{i}=\left(E_{i}-E_{g}\right)/\left(E_{\max}-E_{g}\right)$, where $E_{g}\left(E_{\max}\right)$ is the eigenenergy of the ground (highest excited) state. Based on this, we divide the eigenvalues into $10$ different energy windows, each containing an equal number of eigenvalues, and average samples and all gaps in each window to obtain an $\left\langle r\right\rangle$, as shown in Fig. 1(b). We observe that as the strength of the quasiperiodic potential $\lambda$ increases, the states in the uppermost window become localized first, similar to the single-particle case. Due to the presence of interactions, the mixing of extended and localized states in the middle region of the spectrum disrupts the original position of the ME. However, the mixing of states within the localized region at the top (or the extended region at the bottom) of the spectrum still preserves their localized (or extended) nature, forming many-body localized (or thermalized) states. In the inset of Fig. 1(b), we fix $\lambda=0.8$. The red and blue lines correspond to the $\left\langle r\right\rangle$ of the lowest energy window and the highest energy window, respectively, which converge to $0.53$ and $0.39$ explainB . This indicates that the states in the highest window become localized while the states in the lowest window remain thermalized, suggesting the existence of a MBME. In the following discussion, we fix $\lambda=0.8$ and one-third filling.
Figure 1: (a) IPR of various eigenstates plotted against their corresponding eigenvalues $E$ and the strength of the quasiperiodic potential of the $t_{1}-t_{2}$ model. Black dashed line indicates the approximate mobility edge obtained from Eq. (2). (b) Variation of $\left\langle r\right\rangle$ with energy density ($\epsilon$) and quasiperiodic potential strength $\lambda$. Here $t_{1}=1$ and $t_{2}=0.2$, $U=0.9$, the system size $L=18$, the number of particles $N=6$, and the number of samples is $20$, with each sample corresponding to an initial phase $\phi$. Inset: Variation of $\left\langle r\right\rangle$ with system size at fixed $\lambda=0.8$ and one-third filling, for the highest (blue) and lowest (red) energy window.
II.2 Dissipation-induced transition between thermal and many-body localized states
@@
where $a=\pm 1$, and $j=1,2,\ldots,L-1$. This form of dissipation can be implemented using cold atoms in optical superlattices BD1 ; BD2 ; BD3 ; BD4 or through arrays of superconducting microwave resonators BD6 . This bond dissipation maintains the particle number but modifies the relative phase between adjacent lattice sites. It synchronizes them from an in-phase (out-of-phase) mode to an out-of-phase (in-phase) mode when $a$ is set to $-1$ ($1$), enabling control over the steady state’s position within the energy spectrum BD5 ; BD7 .
@@
When dissipation is present, we note that even if the steady state of the system resides in a localized region under the basis of the Hamiltonian, the system’s dynamics are not necessarily localized. This is because the bond dissipation described by Eq. (3) in our study expands to include terms related to dephasing dissipation $c^{\dagger}_{j}c_{j}$, which which can destroy localization by disrupting coherence. Therefore, we also need to remove dissipation after the system reaches a steady state. Dephasing dissipation destroys localization, but once dissipation is removed after the system reaches a steady state, coherence is restored, and the system recovers its localized state. However, the dissipation we introduced modifies the MBL or thermal states through a different mechanism, namely by selecting certain states. Once the system reaches a steady state, this state selection has already been completed. Upon removing dissipation, the effects induced by the bond dissipation do not vanish. This can be observed through the evolution of the density matrix after dissipation is removed. After abruptly removing the dissipation, we observe the evolution of the density matrix: $\rho(t)=e^{iHt}\rho e^{-iHt}=\sum_{mn}e^{i(E_{m}-E_{n})t}\rho_{mn}|\psi_{m}\rangle\langle\psi_{n}|$, where $E_{m}$ and $E_{n}$ are the eigenvalues corresponding to the many-body eigenstates $|\psi_{m}\rangle$ and $|\psi_{n}\rangle$. We observe that the diagonal elements of $\rho$ (where $E_{m}=E_{n}$) do not change with time, while the off-diagonal elements fluctuate over time. However, with longer observation times, we need to examine the average effects of the dynamics over this interval, leading to the vanishing of the effects caused by the off-diagonal elements explainoff . Therefore, when the steady state primarily occupies the localized (thermal) region, removing dissipation reveals behavior characteristic of the localized (thermal) region. A superposition of localized (thermal) states remains localized (thermal). Based on this observation, we can conclude that dissipation serves as an intermediate process capable of inducing a transition between thermalization and MBL states.
Changes between thermalization and MBL induced by dissipation will inevitably cause changes in transport properties review2 ; Basko2006 ; Gornyi2005 ; Varma2016 , and these transport properties can be probed by experimental platforms such as ultracold atoms transportExp1 ; transportExp2 ; transportExp3 ; transportExp4 ; transportExp5 . By using these properties to distinguish between thermalization and MBL behaviors, we can experimentally verify that dissipation can induce transitions between thermal and localized states. We assume the system reaches a steady state at the time $t_{0}$, after which dissipation is removed. We mainly focus on the transport properties after this point. The system then returns to a non-equilibrium state, and we study its response to a probed electric field. The change in current $\delta I$, for a weak probed electric field, can be derived from linear response theory.
@@
$\sigma(t,t^{{}^{\prime}})=\frac{i}{N}\theta(t-t^{{}^{\prime}})Tr[\rho(t_{0})[I(t),B(t^{{}^{\prime}})]],$ (6)
where the position operator $B=\sum_{j}jc_{j}^{\dagger}c_{j}$ and the current operator $I=\frac{dB}{dt}=-i[B,H]$. The current change in the thermal region is expected to show large fluctuations because it is significantly affected by the electric field, whereas in the MBL state, the current change should exhibit smaller fluctuations due to minimal influence from the electric field (see Appendix C). In Fig. 3, using the same parameters as in Fig. 2, we show $\delta I$ for the cases $a=1$ (blue line) and $a=-1$ (orange line). When $a=1$, as discussed earlier, the steady state primarily occupies the thermalization region near the ground state. After removing dissipation, we observe that the current change consistently exhibits larger oscillations. In contrast, when $a=-1$ (the orange line in Fig. 3), the system’s steady state mainly occupies the many-body localized region near the highest state. After removing dissipation, we see that the current change oscillates with a smaller amplitude and evolves to near zero at a faster rate, exhibiting relatively localized properties.
@@
Figure 4: (a) the distribution of $\delta n$ at each lattice site after adding a particle to a system with $L=144$ and $N=32$. (b) The variation of SIPR with system size. The orange and blue dots represent the ground state and the highest excited state, respectively. Here, we always retain $n_{s}=40$ states when truncating the reduced density matrix and perform $20$ sweeps in the DMRG calculation.
II.3 Discussion on the effects of system size
As discussed earlier, in small systems, there exists a MBME, and dissipation can induce transitions between thermal and MBL states. However, when the system size increases, the existence of a MBME becomes a controversial issue. If there is no MBME, then the dissipation-induced transitions we discussed here would not exist. Inspired by recent experiments DeteME1 ; DeteME2 that probe the ME by measuring the localized and extended properties of the ground state and the highest excited state, we use the density matrix renormalization group (DMRG) method to study the thermalization and localization properties of the ground state and the highest excited state in large systems. To characterize the localization properties of these two states, we consider the particle insertion response, defined as the difference in distribution after introducing an additional particle into the many-body system, i.e., $\delta n_{j}=\rho_{N+1}\left(j\right)-\rho_{N}\left(j\right)$, where $\rho_{N}\left(j\right)=\langle\psi^{N}_{g(e)}|n_{j}|\psi^{N}_{g(e)}\rangle$ and $|\psi^{N}_{g(e)}\rangle$ is the ground (highest excited) state of the system with $N$ particles. Such a quantity has been used to characterize the topologically localized edge excitations Hu1 ; Hu2 and the thermal-MBL transition of the ground state in many-body systems WangEPJB . Similarly, we can also define the IPR of single-particle (SIPR) density change as